Link BD is made of brass (E 105 GPa) and has a cross-sectional area of 240 mm2. Link CE is made of aluminium (E 72 GPa) and has a cross-sectional area of 300 mm2. Knowing that they support rigid member ABC, determine the maximum force P that can be applied vertically at point A if the deflection of A is not to exceed 0.36 mm. (Round the final answer to two decimal places.) 995 mm A 150 mm E 225 mm 125 mm kN. The maximum force P that can be applied vertically at point A is Previous attempt
Each of the rods BD and CE is made of brass (E 105 GPa) and has a cross- sectional area of 200 mm2. Determine the deflection of end A of the rigid member ABC caused by the load P 6 kN. (Round the final answer to three decimal places.) E 225 m S30 mn 100 m The deflection of end A is mm (Click to select) Previous attempt
Answer #1

free bad AC Saluton l A . 225 FBD O 0 35P FBD156 P ( 0 125P-0 225FE = 0 E= 0.56P SB Fe LeD(15P) (0.225) Aeb- Epo (osxi(24xt69S8+Se Lec S lope = 0 225 78-87 XloP -3-8X69P(o125)78-87 xl6-0) + 23 7y8 XloP Apelyy dapbament imit, 27.748x /oP 036 Xlo m= SASCE Smm oomm SaD uaotry uaiyino gepfedia ScE SBD +5 loo SeD 0.75 6cE tlat Sabetitadingy For equaton () ard (2) m (3) and Lep-SA 0-689 SBD (sL+ 05) SA =5.4mm